BRC 3000 Chassis MKII
Model: HMV 2703
Original List Price : £251.00
CRT: A49 120X
This set has seen many hours and frankly has been subject of poor repairs/ maintenance over the years. Added to that it was stored badly for 20+ years before it came into my possession. The cabinet should respond to some oil and be fairly presentable, the electronics within will be another matter entirely. This is what I intend to cover now.
The HMV 2703
PSU and LTB are Junk, they have suffered over the years and have been badly maintained. In fact they have to be the worst I have seen. yet. I could spend hours repairing this to some sort of working condition but frankly they are never going to be reliable or trustworthy. Therefore I have decided these two modules will be replaced with refurbished units from my stores.
The CRT was know to be poor and this was confirmed (Photo 1) with the B&K CRT tester. Red and Green responded to a clean and balance Blue did not (Photo2 ). Blue was given a bop and now reads as good as Red and Green (Photo 3).
The rest of the panels are in reasonable to good condition and should prove to be fine. As an aside I found some more bodging, the main chassis trip had a single strand of wire strapped around the terminals. I removed the trip and fitted and brand new on from stock.
New PSU and LTB tested
All Panels Removed for inspection
As mentioned above the original PSU module looked rough and I mean rough so I thought I would try the spare PSU as shown above. Once installed that gave me a raster albeit a reduced one. For some reason the 60V rail appears to be dragged down and only 8kV is obtained.
So it looks like my spare PSU module has a fault, a 60V rail drain despite appearing and testing OK on the Thorn Test Jig
I decided to test the original PSU and remarkably despite its awful condition appears to be OK. I reinstalled into the set and powered on. Up came a full raster and around 22-23kV of EHT, just shows looks can be deceiving.
I did some major surgery to the PCB which was cracked in two places. This involved drilling a few holes and stitching with strong wire twists, finally soldering these points. Not pretty but an effective and strong mechanical repair. It will do until I can figure out why the spare PSU has the drain.
With the original PSU installed and working, this was what I obtained. with a 625 signal fed to the set, a rather poorly signal could be just about obtained.
I had my suspicion with regards to the I.F. panel so for a quick test a known good panel was installed. This proved my assumption correct, the old I.F. indeed had a fault as a nicely locked test card F was presented.
From this point I could see colour running through at the top of the test card suggesting the reference oscillator was off. At this point the PSU died. Voltage checks showed I had the 240V rail, W605 30V etc was present. I thought here we go again BRC 3000 PSU’s Turns out at this precise time R607 which is a section of the main dropper went high and I mean high, from 100R to 10M. Replaced the dropper, powered on still nothing. Found R645 on the base of VT601 had also decided to go o/c. Replaced and this time I was back in business.
Next job will be to put the old I.F. panel back in and locate the source of the I.F. fault. Then move on to the reference oscillator fault on the chroma panel.
Looks like I’m not getting away with other fault finding, a squealing and squeegeeing sound started emanating from the top of the chassis. Immediate thoughts were “oh god here goes the line output transformer”. EHT was fluctuating and the raster or Test card F had wriggly verticals.
Powered down and installed another LTB to confirm if this was indeed the trouble, low and behold the squealing was still there. Two fault line time-bases? I think not. Oh deep joy its going to be the PSU, sticking my ear a tad closer to the PSU module I could hear the ringing/squealing sound much more prominent and I think I located the source.
Watching the EHT stable then suddenly fluctuate and the location of the straining sounds, I believe the PSU’s Chopper Drive transformer T602 has decided to put its boots up in the air and is failing, Fan-bloomin-tastic!!! I suppose with the state of this set it not to be unexpected and I have a feeling there’s a whole load more fun and games to be had yet.
Found a new Chopper tx (see above) later I will perform some additional PSU surgery and transplant it in.
Replaced the Chopper TX, the old looked rusty and I thought was the cause but no. With a new T601 installed I fired the PSU up on the Thorn test rig and the ringing could be heard although not as dramatic as when in the set. I put this down to not being under the same high load condition it would experience in the TV. I guess I’m looking at component/s in the drive circuit that’s causing this problem.
Whilst studying the cct for likely candidates for the ringing Chopper tx, I notice it had stopped. A glance over that the jig showed the 58-65 rail had died. Time for the scope to see what was happening around the mono-stable multi-vibrator VT603/VT606 and the chopper drive VT605
|VT604||.340||.019||335 (three hundred thirty five)|
A good friend and ex D|E|R /Thorn TV engineer made the following comments
“This does remind me of the problems you had with your 8749 Chris. The waveforms on VT603 base and collector look good but something is certainly awry on VT605. If you isolate the base on VT605 does the collector rise to around 12V? if so something to look at would be around VT606.
Now this loss of the 58-65 rail happened after you changed the chopper drive transformer. These are very reliable but is the one you fitted a known good one?
Remove VT605 completely and measure the collector pad. If it goes up to around 25-30V then VT605 is probably toast. If its still low and R607 overheating then check C608, C633.
For peace of mind I put the old Chopper TX back in the module, still no 58-65V rail. Then isolated Base of VT605 where its collector sits at 3.4V
|Capacitors||C604, C610, C611, C617, C616, C618, C620, C623, C627, C631|
|Resistors||R605, R606, R607, R612, R616, R619, R620,|
Tested VT605 and there is a short between Emitter & Collector, once replaced the 58-65 rail was restored.
Back to this TV repair
After repairing the fault on this HMV’s PSU and then getting engrossed and side-tracked in three other Thorn PSU module repairs, I’m back.
This morning I removed the chroma board for further inspection. As you all will be aware by now this set prior to my ownership, had been poorly stored. As with the other sets from that batch corrosion issues are going to play a big part in the faults this set presents, let alone the usual age related and normal stock faults.
Previously I had discovered W322 (OA91) the pulse clipper o/c so it was replaced. Today on further inspection all the presets look to have lead-out corrosion, if this is causing faults it remains to be seen. However one in particular R354 which controls the pulse width, just fell apart when tapped.
This preset is normally left fully anti-clockwise if it is too far clockwise, the effect will be colour lost could this be thecandidate for my lack of colour. L305 the chrominance peaking coil looks bad but one thing at a time, I will fit a new 47K skeleton and see what we get from there.
The Power-up after PSU and R354 Replacement
The repaired PSU was installed with a rough EHT set-up via the chopper fuse setting it for 60V. R354 (Pulse width) 47K skeleton replaced. The Chroma board was cleaned, rinsed and left in the sun to dry. A few hours later we are ready for power up.
Cleaned up Chroma
Thankfully no dramatics from the the PSU and EHT sits at around 23kV, the nasty squeeging from the line timebase eliminated. Unfortunately my assumption that the open R354 pulse width was likely the cause of no colour, proved to be wrong, I still have no colour.
The resultant B&W test card is not half bad.
Next step is to disable the CK (colour killer) to ascertain what the colour fault actually is. It could be simple as ref osc not locked needing adjustment. Could be the crystal or…… well we shall see. To disable the CK you connect and 82K resistor from the junction of C323/C324 to chassis.
Still no colour when overriding the colour killer. Reading up this suggests either the chrominance signal or the reference sub-carrier is missing. I can hear shouts of “change the crystal”, yes I’ve heard that before and it is possible but on all occasions when I was told to do that, it turned out not to be the crystal. Sods law it will be now! Anyway before that I’m going to get the scope on it.
Scoped the chroma board, its shut down no gated burst, no ½ line frequency component, no ref. This one is going to take a while to diagnose especially with all the rusted component leads, locfits and L305 with a bad dose of green spot or splodge in its case.
Before I dive into the fault finding I’m desperate to see if the CRT is likely to give a good account of itself so I grabbed a spare chroma of unknown state from the stock pile. Packed with callin caps it worked first time and the CRT does look like it will give a reasonable picture.
OK no cheating allowed, remove spare Chroma, deep breath, back in with the HMV chroma panel and head down for some fault finding.
Update before bed
With the original chroma board back in I have not colour, replaced the 4.43 crystal and I now have barber-pole, unlocked colour. The oscillator is now running and I need to tweak it to lock, that’s for tomorrow.
Update: Decoder diag day.
Starting from where I left off last night I have an unlocked reference oscillator (above) which is due to the reference oscillator not being locked to the transmitted burst.
Therefore the first thing I looked for with the scope was the burst waveform at the collector of VT302, it was not present. This might be due to the gated burst amplifier VT301 not working, so I checked for the gating pulse waveforms in and out of the polarity splitter VT308, again nothing.
Now VT308 provides the anti-phase pulses to drive the burst blanking diodes and to supply the burst gating, so if this transistor has failed there’s no burst, no ident and no colour. This is the one I will start at first. I will take some voltage readings on the other transistors then its out with the board, remove and test VT308 not forgetting whilst its out VT301, VT302, W302 and W303.
It was indeed VT308 that had failed
With VT308 replaced I now I had burst at the collector of VT302 and the gating pulse going in and out of the polarity splitter
Took a voltage map of all the silicon on the decoder, the problem looks to be in or around VT303 the DC amp. I should only have 1.1v on its emitter an 4v on the collector. As can be seen below mine are sky high.
Reading up on the 3K chrominance circuit, the amplified burst is compared with the reference oscillator signal in phase discriminator L301-W302-W303 and the appropriate correction is applied to oscillator via the DC Amplifier VT303. With VT303 not working this correction will not be occurring, so is this the cause of the inability to lock.
|VT301||Burst Gate & Amp||1.7||5.6||29.1|
|VT302||Burst Gate & Amp||29.1||29.2||.013|
|VT303||D.C. Amp||14.9||24.1 (1.1)||29.1 (4)|
|VT307||Switched Amp||27.6||28.2||13.2 (17)|
|VT310||Delay Line Driver||13.7||13.1||25.7|
First I removed and test VT303, it was OK. R315 which was erratic so it was replaced. V302 and W303 were tested and fell apart when probed despite looking OK, more rust issues. R312 was also replaced as the fell apart. These changes now allowed me to get a locked colour albeit incorrect colour.
Following a quick decoder set-up using a DMM I obtained correct colour bar. Feeding in Test Card F and a pleasant result. I will do a full and proper decoder set-up with the scope a bit later.
HMV Decoder: Starting to give up the fight
Its starting to look fairly good now.
One thing that needed addressing was the crooked scan coil alignment, they were unlocked then corrected.A convergence was then required.
Here its is after, It’ll likely improve further once full installation and set up as per the service manual, which is next.
Installation Set Up
- Set AFC
- Set AGC
- Set Porch Bias
- Set Clamp Pulse
- Set Clamp Pulse Phase
- Clamp Pulse Amplitude
- Grey-Scale Tracking
- Set ACC
- CRT Grid Bias
- SET EHT
- Preset Brightness
- Beam Limiter
A whole lot better after another grey-scaling and convergence.
Cabinet all cleaned a refinished with new HMV badge fitted.